OpenStudy (help!!!!):

Matt's Oak Superstore has exactly three times as mant large oak desks as small oak desks in its inventory. If the store only sells these two types of desks, which could be the total number of desks in stock?

OpenStudy (aripotta):

are there answer choices?

OpenStudy (help!!!!):

yeah

OpenStudy (aripotta):

can you post em? :l

OpenStudy (help!!!!):

give me a sec

OpenStudy (help!!!!):

a. 10 b.13 c.16 d.18 e.25

OpenStudy (aripotta):

c

OpenStudy (help!!!!):

why c?

OpenStudy (aripotta):

3x + x = ? 4x = ? 4x = 10 x = 2.5 (can't be it) 4x = 13 x = 3.25 (can't be it) 4x = 16 x = 4 (the only choice that CAN be it) 4x = 18 x = 4.5 (can't be it) 4x = 25 x = 6.25 ( can't be it)

OpenStudy (aripotta):

does that make sense?

OpenStudy (help!!!!):

si

OpenStudy (help!!!!):

but can you fan me?

OpenStudy (aripotta):

uh sure. fan for fan?

OpenStudy (help!!!!):

kk